# A Survey on Transformer Compression

### 05 Feb 2024 v1

@@ Abstract
Large models based on the Transformer architecture play increasingly vital roles in artificial intelligence, particularly within the realms of natural language processing (NLP) and computer vision (CV). Model compression methods reduce their memory and computational cost, which is a necessary step to implement the transformer models on practical devices. Given the unique architecture of transformer, featuring alternative attention and Feedforward Neural Network (FFN) modules, specific compression techniques are required. The efficiency of these compression methods is also paramount, as it is usually impractical to retrain large models on the entire training dataset. This survey provides a comprehensive review of recent compression methods, with a specific focus on their application to transformer models. The compression methods are primarily categorized into pruning, quantization, knowledge distillation, and efficient architecture design. In each category, we discuss compression methods for both CV and NLP tasks, highlighting common underlying principles. At last, we delve into the relation between various compression methods, and discuss the further directions in this domain.
@@
Model Compression, Transformer, Large Models, Computer Vision, Natural Language Processing
@@ 1 Introduction
Deep neural networks have become indispensable in numerous artificial intelligence applications, whose architectures encompass diverse formulations, such as multilayer perceptron (MLP), convolutional neural network (CNN), recurrent neural network (RNN), long short-term memory (LSTM), Transformers, etc. In recent times, transformer-based models have emerged as the prevailing choice across various domains, including both natural language processing (NLP) and computer vision (CV) domains. Considering their strong scaling ability, most of the large models with over billions of parameters are based on the transformer architecture, which are considered as foundational elements for general artificial intelligence (AGI) [1, 2, 3, 4, 5, 6].
@@
While large models have demonstrated significant capabilities, their exceptionally vast sizes pose challenges for practical development. For instance, the GPT-3 model has 175 billion parameters and demands approximately about 350GB memory model storage (float16). The sheer volume of parameters and the associated computational expenses necessitate devices with exceedingly high memory and computational capabilities. Directly deploying such models will incur substantial resource costs and contributes significantly to carbon dioxide emissions. Moreover, on edge devices like mobile phones, the development of these models becomes impractical due to their limited storage and computing resource.
@@
Figure 1: Transformer-based models have emerged as the predominant architectures in both natural language processing (NLP) and computer vision (CV) domains, garnering a surge in publications. As these models tend to possess substantial dimensions, the imperative arises to compress their parameters and streamline computational redundancies. This compression is essential for facilitating efficient implementation on practical platforms, ensuring the feasibility of deploying Transformer models in real-world applications.
@@
Model compression is an effective strategy for mitigating the development costs associated with transformer models. This approach, grounded in the principle of reducing redundancy, encompasses various categories, including pruning, quantization, knowledge distillation, efficient architecture design, etc. Network pruning directly removes redundant components, such as blocks, attention heads, FFN layers or individual parameters. Diverse sub-models can be derived by employing different pruning granularity and pruning criteria. Quantization reduces the development cost by representing model weights and intermediate features with lower bits. For example, When quantizing a full-precision model (float32) into 8-bit integers, the memory cost can be reduced by a factor of four. According the computational process, it can be divided into post-training quantization(PTQ) or quantization-aware training (QAT), in which the former only incurs limited training costs and more efficient for large models. Knowledge distillation serves as a training strategy, which transfers knowledge from a large model (teacher) to a smaller model (student). The student mimics the behavior of the teacher by emulating the model’s output and intermediate features. Notably, for advanced models like GPT-4, accessible only through APIs, their generated instructions and explanations can also guide the learning of the student model [7, 8].In addition to obtaining models from predefined large models, some methods yield efficient architectures by directly reducing the computational complexity of attention modules or FFN modules. Combining different methods enables extreme compression. For instance, Han et al. [9] combined network pruning, quantization, and Huffman coding, achieving an impressive 49$\times$ compression rate on a conventional VGGNet [10].
@@
Regarding transformer models, their compression strategies exhibit distinct characteristics. Different from other architectures like CNN or RNN, the transformer features a unique design with alternative attention and FFN modules. The former captures the global information by calculating the attention map over different tokens while the other extracts information from each token respectively. This specific architecture can inspire tailored compression strategy for optimal compression rates. Besides, the efficiency of compression method becomes especially important for such large models. Due to the high computational cost of large model, it is usually unaffordable to retrain the whole model on the original training set. Some training-efficient methods like post-training compression are more preferable.
@@
In this survey, we aim to comprehensively investigate how to compress these transformer models (Figure 1), and the methods are categorized by quantization, knowledge distillation, pruning, efficient architecture design, etc. In each category, we investigate the compression methods for NLP and CV domains, respectively. Table I summarizes the main compression categories and lists representative method suitable for large transformer models. Though NLP and CV are usually treated as very different domains, we observe that their models compression methods actually share the similar principles. Finally, we discuss the relationship between different compression methods and post some future research directions.
@@
The rest of the paper is organized as follows. Section 2 introduces the fundamental concept of transformers. Following this, Section 3 provides an in-depth discussion on compression methods that preserve the architecture, encompassing quantization and knowledge distillation—techniques that maintain the model’s architecture. Section 4 delves into further exploration of architecture-preserving compression, including pruning and efficient architecture design. Additional transformer compression methods are explored in Section 5. Finally, Section 6 concludes the compression method and discuss future research directions.
@@ 2 Concept of Transformer
where $Q$, $K$, $V$ are query, key and value matrix, respectively. $d$ is the feature’s dimension. The multi-head attention jointly extracts information from diverse subspaces, which is the concatenation of different heads, i.e.,
@@
$W^{Q}$, $W^{K}$, $W^{V}$, $W^{O}$ are the corresponding parameter matrices. The FFN module transforms features from each token independently. It is usually constructed by stacking two FC layers with activation functions, i.e.,
@@
where $x$ is input feature, $\phi$ is activation function (e.g., GeLU). $W_{1}$, $W_{2}$, $b_{1}$, $b_{2}$ are the weight and bias parameters in FC layers. The MHA and FFN module are stacked alternatively to construct the whole model.
The Transformer architecture has strong scaling ability to construct extremely large models with several billion or trillion parameters. It supports most of the prevalent large models in NLP, CV and multiple modality domains. For example, the well-known large language models (e.g., GPT-series [4, 2], LLAMA [1],Pangu [5, 6]) are its decoder-only variants. By simply splitting an image into multiple patches, it can be used to tackle vision tasks [12, 13, 14]. The multiple model like CLIP [15], BLIP [16], LLaVA [17] also use transformer as the backbones.
@@
Sheared LLaMA [30] Structured, Pre-defined tagert, dyanmic data loading NeurIPS 2023
@@ 3.1.1 Overview of Quantization
Quantization is a crucial step for deploying transformers on various devices, especially GPUs and NPUs, that design specialized circuits for low-precision arithmetic. During the quantization process as Equation 4, a floating-point tensor $x$ is converted to the integer one $x_{int}$ with corresponding quantization parameters (scale factor $s$ and zero point $z$), then the integer tensor $x_{int}$ could be quantized back to floating-point $x_{quant}$ but causes some precision error compared with original $x$, i.e.,
@@
where $s_{w},s_{e},z_{e}$ are quantization parameters of weights and input embedding, and $e_{int}$ and $\textrm{w}_{int}$ are the integer input and weights, which calculated by Equation 4. $C$ could be pre-computed with $s_{e},z_{e},s_{e}$ and $\textrm{w}_{int}$. Thus the floating-point multiplication could be accelerate with efficient integer multiplication in the inference. To minimize the performance degradation of quantized models, different optimization methods have been proposed and they can be divided into two categories: (1) Post-training quantization (PTQ) [21, 41, 22, 42, 43, 44, 45] , mainly focuses on optimizing the quantization parameters of weights and activations with a few unlabeled calibration data, some latest methods also explore adaptive rounding for weight quantization. (2) Quantization-aware training (QAT) [46, 47, 48, 49, 50, 51, 23, 52, 53, 54, 55, 56], inserts the quantization nodes into networks and conducts training with complete training data, where all the weights and quantization parameters are optimized together. In this section, we systematically introduce the research of model quantization on transformer-based vision models and large language models, as shown in Figure 2.
Figure 2: The overview of quantization for transformers. The top demonstrates the different problems that addressed in existing works for computer vision and natural language processing, and the bottom shows a normal INT8 inference process of a standard transformer block.
@@ 3.1.2 Quantization for Transformer-Based Large Language Models
Before 2023, the study of quantization [57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67] on transformer-based NLP almost focuses on BERT architecture. With the popularity of pretrained large language models [68, 1, 2], researchers [69, 70, 71] explore how to quantize the transformer with Billion-level parameters and explores more efficient quantization schemes with limited data and computational overhead.
@@
Post-training Quantization. Based on the analysis of outliers for quantized transformers, Outlier Suppression [72] proposes to migrate the gamma of LayerNorm to the next module for that the gamma amplifies the outliers in the output and causes large quantization error, and then clips the tokens in a coarse-to-fine procedure. MREM [59] focuses on reducing the computational cost of quantization, presents to minimize the output quantization error for all modules in parallel manner. Zeroquant [71] finds that the performance degradation comes from the different dynamic range of tokens and weights, proposes to adopt group-wise quantization for weights and token-wise quantization for activations, and utilizes knowledge distill layer by layer. For large transformers with billion-level parameters, the outlier is still the main reason that resulting large accuracy degradation for quantized models. To address that, LLM.int8() [69] represents the activations and outliers of weights with 16-bit and performs 8-bit vector-wise quantization for weight tensor, but the acceleration is limited and even worse due to the irregular quantization granularity [18]. GPTQ [73] also only quantizes weight parameters as LLM.int8(), but adopts unify quantization strategy. GPTQ quantizes the weights based on approximate second-order information, which could get much accurate quantized weight in a few hours. AWQ [74] proposes to search the optimal per-channel scale factors by observing the distribution of activations instead of weights, thus the LLMs could remain the generalist abilities for different domains and modalities. Outlier Suppression+ [75] explores more accurate outliers suppression scheme with channel-wise shifting and scale, which helps align the range of different activation channels and scale down the outliers. The shifting and scale factors could also be merged with other weight parameters. Similarly, SmoothQuant [18] and QLLM [76] propose the mathematically equivalent per-channel scaling transformation that migrates the quantization difficulty from activations to weights. To further improve the performance of quantized LLMs, QLLM also learns a low-rank parameters by minimizing the reconstruction error of the outputs between floating-point and quantized LLMs with limited calibration data. Similarly, RPTQ [77] utilizes a reorder-based scheme that rearranging the channels of activations with similar range and quantizes them in cluster, and then migrates the scale into LayerNorm and weights of linear layers without extra computational overhead in the inference. Based on scale migration, OmniQuant [19] further proposes a learnable PTQ methods module by module, where the weight clipping parameters and transformation scale of activations are optimized with gradient decent. Like GPTQ, SqueezeLLM [78] also focuses on quantizing weights for the memory bandwidth and proposes to search the optimal bit based on second-order information. Besides, SqueezeLLM does not suppression the outliers and sensitive weight values but store them in a efficient sparse format for more accurate quantized LLMs. Different from the previous methods, SignRound [79] proposes to optimize quantized LLMs from the perspective of adaptive rounding. SignRound designs block-wise tuning using signed gradient descent to learn the weight rounding, which greatly help the output reconstruction of each block.
@@
Quantization-aware training. Q-BERT [57] is the early work that conducts quantization-aware training on transformer-based architecture for natural language processing. Inspired by HAWQ [82], Q-BERT searches mixed-quantization setting based on the second order Hessian information and adopts group-wise quantization to partition parameter matrix into multiply groups to reduce accuracy degradation. I-BERT [62] not only quantizes the linear and self-attention layer, but also designs an integer-only inference scheme for the nonlinear operations (GELU, SoftMax and LayerNorm) in transformers, which also inspires the FQ-ViT and I-ViT for the quantization of vision transformer. Specifically, I-BERT proposes to use the second-order polynomial approximate for GELU and exponential function of SoftMax, and calculate the standard deviation of LayerNorm based on Newton’s Method, thus I-BERT could achieve faster inference compared to the baseline and normal quantization methods. To compensate for the disadvantages of the hand-crafted heuristics in Q-BERT, AQ-BERT [66] proposes an automatic mixed-precision quantization scheme to learn the bit-width and parameters of each layer simultaneously, which is inspired from differentiable network architecture searching. AQ-BERT achieves better results than Q-BERT and more suitable for resource-limited devices. Except the QAT schemes for BERT, there are only a few papers that conducts quantization-aware training on large language models for the huge training overhead. PEQA [81] and QLoRA [20] explore parameter-efficient fine-tuning techniques to train the quantized LLMs, the former only optimizes the scale factors while freezing the quantized weights, and the later adopts 4-bit NormalFloat, double quantization and a paged optimizers for normally distributed weights and reduce memory footprint during training. LLM-QAT [80] proposes a data-free quantization-aware training methods, where the training data are generated from the original pretrained large language model. LLM-QAT not only quantizes all the linear layers and self-attention, but also quantizes the KV cache with cross-entropy based logits distillation.

### 07 Apr 2024 v2

@@ Abstract
Transformer plays a vital role in the realms of natural language processing (NLP) and computer vision (CV), specially for constructing large language models (LLM) and large vision models (LVM). Model compression methods reduce the memory and computational cost of Transformer, which is a necessary step to implement large language/vision models on practical devices. Given the unique architecture of Transformer, featuring alternative attention and feedforward neural network (FFN) modules, specific compression techniques are usually required. The efficiency of these compression methods is also paramount, as retraining large models on the entire training dataset is usually impractical. This survey provides a comprehensive review of recent compression methods, with a specific focus on their application to Transformer-based models. The compression methods are primarily categorized into pruning, quantization, knowledge distillation, and efficient architecture design (Mamba, RetNet, RWKV, etc). In each category, we discuss compression methods for both language and vision tasks, highlighting common underlying principles. Finally, we delve into the relation between various compression methods, and discuss further directions in this domain.
@@
Model Compression, Transformer, Large Language Model, Large Vision Model, LLM
@@ 1 Introduction
Deep neural networks have become indispensable in numerous artificial intelligence applications, with architectures encompassing diverse formulations, such as multilayer perceptron (MLP), convolutional neural network (CNN), recurrent neural network (RNN), long short-term memory (LSTM), Transformers, etc. In recent times, Transformer-based models have emerged as the prevailing choice across various domains, including both natural language processing (NLP) and computer vision (CV) domains. Considering their strong scalability, most of the large models with over billions of parameters are based on the Transformer architecture, which are considered as foundational elements for artificial general intelligence (AGI) [1, 2, 3, 4, 5, 6].
@@
While large models have demonstrated significant capabilities, their exceptionally vast sizes pose challenges for practical development. For instance, the GPT-3 model has 175 billion parameters and demands about 350GB memory model storage (float16). The sheer volume of parameters and the associated computational expenses necessitate devices with exceedingly high memory and computational capabilities. Directly deploying such models will incur substantial resource costs and contributes significantly to carbon dioxide emissions. Moreover, on edge devices like mobile phones, the development of these models becomes impractical due to the limited storage and computing resources of such devices.
@@
Figure 1: Transformer-based models have emerged as the predominant architectures in both natural language processing (NLP) and computer vision (CV) domains, resulting in a surge in publications. As these models tend to possess substantial dimensions, it becomes imperative to compress their parameters and streamline computational redundancies. This compression is essential for facilitating efficient implementation on practical platforms, ensuring the feasibility of deploying Transformer models in real-world applications.
@@
Model compression is an effective strategy for mitigating the development costs associated with Transformer models. This approach, grounded in the principle of reducing redundancy, encompasses various categories, including pruning, quantization, knowledge distillation, efficient architecture design, etc. Network pruning directly removes redundant components, such as blocks, attention heads, FFN layers, and individual parameters. Diverse sub-models can be derived by employing different pruning granularity and pruning criteria. Quantization reduces the development cost by representing model weights and intermediate features with lower bits. For example, when quantizing a full-precision model (float32) into 8-bit integers, the memory cost can be reduced by a factor of four. According the computational process, it can be divided into post-training quantization(PTQ) or quantization-aware training (QAT), in which the former incurs only limited training costs and is more efficient for large models. Knowledge distillation serves as a training strategy, and transfers knowledge from a large model (teacher) to a smaller model (student). The student mimics the behavior of the teacher by emulating the model’s output and intermediate features. Notably, for advanced models like GPT-4, accessible only through APIs, their generated instructions and explanations can also guide the learning of the student model [7, 8].In addition to obtaining models from predefined large models, some methods yield efficient architectures by directly reducing the computational complexity of attention modules or FFN modules. Combining different methods enables extreme compression. For instance, Han et al. [9] combined network pruning, quantization, and Huffman coding to achieve an impressive 49$\times$ compression rate on a conventional VGGNet [10].
@@
Regarding Transformer models, their compression strategies exhibit distinct characteristics. Unlike other architectures such as CNN or RNN, the Transformer features a unique design with alternative attention and FFN modules. The former captures the global information by calculating the attention map over different tokens while the latter extracts information from each token respectively. This specific architecture can enable a tailored compression strategy for optimal compression rates. What’s more, the efficiency of compression method becomes especially important for such large models. Due to the high computational cost of large model, it is usually unaffordable to retrain the whole model on the original training set. Some training-efficient methods like post-training compression are preferable.
@@
In this survey, we aim to comprehensively investigate how to compress these Transformer models (Figure 1), and categorize the methods by quantization, knowledge distillation, pruning, efficient architecture design, etc. In each category, we investigate the compression methods for NLP and CV domains, respectively. Table I summarizes the main compression categories and lists representative method suitable for large Transformer models. Though NLP and CV are usually treated as very different domains, we observe that their models compression methods actually share the similar principles. Finally, we discuss the relationship between different compression methods and outline some future research directions.
@@
The rest of the paper is organized as follows. Section 2 introduces the fundamental concept of Transformers. Following this, Section 3 provides an in-depth discussion on compression methods that preserve the architecture, encompassing quantization and knowledge distillation—techniques that maintain the model’s architecture. Section 4 delves further into architecture-preserving compression, including pruning and efficient architecture design. Additional Transformer compression methods are explored in Section 5. Finally, Section 6 draws conclusions on the compression methods and discusses future research directions.
@@ 2 Concept of Transformer
where $Q$, $K$, $V$ are query, key and value matrix, respectively. $d$ is the feature’s dimension. The multi-head attention jointly extracts information from diverse subspaces, which is the concatenation of different heads,
@@
$W^{Q}$, $W^{K}$, $W^{V}$, $W^{O}$ are the corresponding parameter matrices. The FFN module transforms features from each token independently. It is usually constructed by stacking two FC layers with activation functions,
@@
where $x$ is input feature, and $\phi$ is activation function (e.g., GELU). $W_{1}$, $W_{2}$, $b_{1}$, $b_{2}$ are the weight and bias parameters in FC layers. The MHA and FFN module are stacked alternatively to construct the whole model.
The Transformer architecture has strong scalability and so can be used to construct extremely large models with several billion or trillion parameters. It supports most of the predominant large models in NLP, CV and multiple modality domains. For example, the well-known large language models (e.g., GPT-series [4, 2], LLaMA [1],Pangu [5, 6]) are its decoder-only variants. By simply splitting an image into multiple patches, it can be used to tackle vision tasks [12, 13, 14]. The multiple model like CLIP [15], BLIP [16], LLaVA [17] also use Transformer as the backbones.
@@
Sheared LLaMA [30] Structured, pre-defined target, dynamic data loading NeurIPS 2023
@@ 3.1.1 Overview of Quantization
Quantization is a crucial step for deploying Transformers on various devices, especially GPUs and NPUs, that have specialized circuits for low-precision arithmetic. During the quantization process as shown in Equation 4, a floating-point tensor $x$ is converted to the integer one $x_{int}$ with corresponding quantization parameters (scale factor $s$ and zero point $z$), then the integer tensor $x_{int}$ could be quantized back to floating-point $x_{quant}$ but causes some precision error compared with the original $x$,
@@
where $s_{w},s_{e},z_{e}$ are quantization parameters of weights and input embedding, and $e_{int}$ and $\textrm{w}_{int}$ are the integer input and weights, which are calculated by Equation 4. $C$ could be pre-computed with $s_{e},z_{e},s_{e}$ and $\textrm{w}_{int}$. Thus the floating-point multiplication could be accelerated with efficient integer multiplication in the inference. To minimize the performance degradation of quantized models, different optimization methods have been proposed and they can be divided into two categories: (1) Post-training quantization (PTQ) [21, 41, 22, 42, 43, 44, 45] , mainly focuses on optimizing the quantization parameters of weights and activations with a few unlabeled calibration data, and some of the latest methods also explore adaptive rounding for weight quantization. (2) Quantization-aware training (QAT) [46, 47, 48, 49, 50, 51, 23, 52, 53, 54, 55, 56], inserts the quantization nodes into networks and conducts training with complete training data, where all the weights and quantization parameters are optimized together. In this section, we systematically introduce the research of model quantization on Transformer-based vision models and large language models, as shown in Figure 2.
Figure 2: The overview of quantization for Transformers. The top demonstrates the different problems that are addressed in existing works for computer vision and natural language processing, and the bottom shows a normal INT8 inference process of a standard Transformer block.
@@ 3.1.2 Quantization for Transformer-Based Large Language Models
Before 2023, the study of quantization [57, 58, 59, 60, 61, 62, 63, 64, 65, 66] on Transformer-based NLP focused almost entirely on the BERT architecture. With the popularity of pretrained large language models [67, 1, 2], researchers [68, 69, 70] stated to explore how to quantize the Transformer with billions of parameters and explores more efficient quantization schemes with limited data and computational overhead.
@@
Post-training quantization. Based on the analysis of outliers for quantized Transformers, Outlier Suppression [71] proposes to migrate the gamma of LayerNorm to the next module for that the gamma amplifies the outliers in the output and causes large quantization error, and then clips the tokens in a coarse-to-fine procedure. MREM [59] focuses on reducing the computational cost of quantization, and minimizes the output quantization error for all modules in parallel manner. Zeroquant [70] finds that the performance degradation is due to the different dynamic range of tokens and weights, proposes to adopt group-wise quantization for weights and token-wise quantization for activations, and utilizes knowledge distillation layer by layer. For large Transformers with billions of parameters, the outlier is still the main reason for large accuracy degradation for quantized models. To address this, LLM.int8() [68] represents the activations and outliers of weights with 16-bit and performs 8-bit vector-wise quantization for weight tensor, but the acceleration is limited and even worsened due to the irregular quantization granularity [18]. GPTQ [72] also only quantizes weight parameters as LLM.int8(), but adopts unified quantization strategy. GPTQ quantizes the weights based on approximate second-order information, thereby getting much more accurate quantized weight in a few hours. AWQ [73] proposes to search for the optimal per-channel scale factors by observing the distribution of activations instead of weights, allowing the LLMs retain the capabilities for different domains and modalities. Outlier Suppression+ [74] explores more accurate outliers suppression schemes with channel-wise shifting and scale, thereby helping align the range of different activation channels and scale down the outliers. The shifting and scale factors could also be merged with other weight parameters. Similarly, SmoothQuant [18] and QLLM [75] propose the mathematically equivalent per-channel scaling transformation that migrates the quantization difficulty from activations to weights. To further improve the performance of quantized LLMs, QLLM also learns low-rank parameters by minimizing the reconstruction error of the outputs between floating-point and quantized LLMs with limited calibration data. Similarly, RPTQ [76] utilizes a reorder-based scheme that rearranges the channels of activations with a similar range and quantizes them in cluster, and then migrates the scale into LayerNorm and weights of linear layers without extra computational overhead in the inference.Based on scale migration, OmniQuant [19] further proposes a learnable PTQ method module by module, where the weight clipping parameters and transformation scale of activations are optimized with gradient descent. To minimize error accumulation in adjacent blocks, CBQ [77] presents a cross-block reconstruction framework, simultaneously learn the rounding matrices of weight and step sizes of weights and activations. The rounding matrices are learned with LoRA technique, which does not bring much extra cost for PTQ. Like GPTQ, SqueezeLLM [78] also focuses on quantizing weights for the memory bandwidth and proposes to search for the optimal bit based on second-order information. Also, SqueezeLLM does not suppress the outliers and sensitive weight values but stores them in an efficient sparse format for more accurate quantized LLMs. Unlike the previous methods, SignRound [79] proposes to optimize quantized LLMs from the perspective of adaptive rounding. SignRound designs block-wise tuning using signed gradient descent to learn the weight rounding, thereby greatly helping the output reconstruction of each block.
@@
Quantization-aware training. Q-BERT [57] is the early work that conducts quantization-aware training on Transformer-based architecture for natural language processing. Inspired by HAWQ [82], Q-BERT searches mixed-quantization settings based on the second order Hessian information and adopts group-wise quantization to partition parameter matrix into multiple groups to reduce accuracy degradation. I-BERT [62] not only quantizes the linear and self-attention layer, but also designs an integer-only inference scheme for the nonlinear operations (GELU, SoftMax and LayerNorm) in Transformers, which also inspires the FQ-ViT and I-ViT for the quantization of vision Transformer. Specifically, I-BERT proposes to use the second-order polynomial approximate for GELU and exponential function of SoftMax, and calculate the standard deviation of LayerNorm based on Newton’s Method. In this way, I-BERT can achieve faster inference compared to the baseline and normal quantization methods. To compensate for the disadvantages of the hand-crafted heuristics in Q-BERT, AQ-BERT [66] proposes an automatic mixed-precision quantization scheme to learn the bit-width and parameters of each layer simultaneously, which is inspired from differentiable network architecture search. AQ-BERT achieves better results than Q-BERT and more suitable for resource-limited devices. Besides the QAT schemes for BERT, there are only a few papers that conducts quantization-aware training on large language models for the huge training overhead. PEQA [81] and QLoRA [20] explore parameter-efficient fine-tuning techniques to train the quantized LLMs, with the former only optimizing the scale factors while freezing the quantized weights, and the latter adopting 4-bit NormalFloat, double quantization, and paged optimizers for normally distributed weights and reduce memory footprint during training. LLM-QAT [80] proposes a data-free quantization-aware training method, where the training data is generated from the original pretrained large language model. LLM-QAT not only quantizes all the linear layers and self-attention, but also quantizes the KV cache with cross-entropy based logits distillation.